OpenStudy (anonymous):

What is the limit as (cosx-1)/(2x) approaches 0?

OpenStudy (lgbasallote):

are you familiar with the l'hospital rule?

OpenStudy (anonymous):

I am not, unfortunately. I have heard of it, though.

OpenStudy (dumbcow):

its one of my fav rules...it involves calculus so if you havn't learned derivatives yet then its of no use to you :|

hartnn (hartnn):

first multiply (cos x + 1) in numerator and denominator and tell me what u get ?

OpenStudy (dumbcow):

he left...yeah that was my thought too but im stuck on next step

hartnn (hartnn):

sorry, my connection kinda lost.... u get -sin^2 x in numerator....

OpenStudy (turingtest):

we all lost our connection

OpenStudy (turingtest):

you guys are over-complicating this first off are you allowed to use l'hospital's rule?

hartnn (hartnn):

actually @SinoatrialMath doesn't know that rule, so i am guiding this sum through traditional way...

OpenStudy (turingtest):

sorry, you were gonna prove lim x->0 (1-cosx)/x algebraically ?

OpenStudy (anonymous):

I'm sorry, I lost connection.

hartnn (hartnn):

yup, not so difficult.

OpenStudy (turingtest):

I have never seen that, please do show @hartnn

OpenStudy (turingtest):

oh I think I see it...

hartnn (hartnn):

ok, step by step.. @SinoatrialMath did u multiply and divide by (1+cos x) ?

OpenStudy (anonymous):

(cos^2x-1)/2x(cosx+1)

hartnn (hartnn):

and whats cos^2 x - 1 ??

OpenStudy (anonymous):

sin^2x

OpenStudy (turingtest):

actually now I don't see it anymore, I wanna see how this turns out too

OpenStudy (anonymous):

(sin^2x)/(-sinx) ?

hartnn (hartnn):

its -sin^2 x so u have \(\huge\frac{-sin^2 x}{2x(cos x+1)}\)

OpenStudy (anonymous):

Ah yes, I'm following.

OpenStudy (turingtest):

for the record, I was just gonna separate the -(1-cosx)/x and 1/2 part... sorry to interrupt

hartnn (hartnn):

now separate ONE of the 2 sin^2 x and write it like this : \(\huge \frac{sin x}{x} \quad \frac{-sin x}{2(cos x +1)}\) ok with this ?

OpenStudy (anonymous):

I understand, yes.

hartnn (hartnn):

now the 1st limit is standard, and its value is 1 for 2nd limit, just put x=0 and tell me what u get ?

OpenStudy (anonymous):

-sin(0) / (2(cos(0)+1)) = 0 / (2(1+1)) = 0 / (2(2)) = 0 / 4 = 0

hartnn (hartnn):

thats right so 1st limit =1 2nd limit =0 overall limit = 1*0 = 0. ok ?

OpenStudy (anonymous):

Wow! Thank you very much! I would have never broken it down like that, taking out one "sinx" as well as the x in the denominator. How did you think of that (taking the x out of the denominator)?

OpenStudy (turingtest):

let me just do mine right quick\[\lim_{x\to0}\frac{\cos x-1}{2x}=-\frac12\lim_{x\to0}\frac{1-\cos x}x\]of which I think the limits\[\lim_{x\to0}\frac{1-\cos x}x=0\]is equally known, so you get\[\lim_{x\to0}\frac{\cos x-1}{2x}=-\frac12\lim_{x\to0}\frac{1-\cos x}x=-\frac12(0)=0\]

OpenStudy (anonymous):

I'm guessing you're quite familiar w/ using that part of the squeeze theorem (lim x-> 0 of sinx/x = 1).

hartnn (hartnn):

its just practice that makes u think in right direction :)

OpenStudy (turingtest):

indeed :)

OpenStudy (anonymous):

Oh that was clever as well, @TuringTest - that was a very nice problem.